# SOLAR 10.7B: Scaling Large Language Models with Simple yet Effective Depth Up-Scaling

### 23 Dec 2023 v1

@@ Abstract
We introduce depth up-scaling (DUS), a novel technique to up-scale base LLMs efficiently and effectively in a simple manner. In contrast to mixture-of-experts (MoE), DUS does not require complex changes to train and inference. Using DUS, we build SOLAR 10.7B, a large language model (LLM) with 10.7 billion parameters, demonstrating superior performance in various natural language processing (NLP) tasks. Comparative evaluations show that SOLAR 10.7B outperforms existing open-source pretrained LLMs, such as Llama 2 and Mistral 7B. We additionally present SOLAR 10.7B-Instruct, a variant fine-tuned for instruction-following capabilities, surpassing Mixtral-8x7B. SOLAR 10.7B is publicly available under the Apache 2.0 license, promoting broad access and application in the LLM field  [^1] https://huggingface.co/upstage/SOLAR-10.7B-v1.0.
@@ 1 Introduction
Figure 1: Depth up-scaling with base models that have 32 layers. We use the Llama2 architecture, but other transformer architectures are compatible with depth up-scaling. We use pretrained weights in 24 of the 32 layers in the base model, which results in the depth up-scaled model having pretrained weights for all of its 48 layers.
The field of natural language processing (NLP) has been significantly transformed by the introduction of large language models (LLMs), which have enhanced our understanding and interaction with human language Zhang et al. (2023a). These advancements, while beneficial, bring challenges such as the increased need to train ever larger models owing to the performance scaling law Kaplan et al. (2020); Hernandez et al. (2021); Anil et al. (2023). To efficiently tackle the above, recent works in scaling language models such as a mixture of experts (MoE) Shazeer et al. (2017); Komatsuzaki et al. (2022) have been proposed. While those approaches are able to efficiently and effectively scale-up LLMs, they often require non-trivial changes to the training and inference framework Gale et al. (2023), which hinders widespread applicability. Effectively and efficiently scaling up LLMs whilst also retaining the simplicity for ease of use is an important problem Alberts et al. (2023); Fraiwan and Khasawneh (2023); Sallam et al. (2023); Bahrini et al. (2023).
Our study introduces depth up-scaling (DUS), an effective and efficient method to up-scale LLMs whilst also remaining strikingly easy to apply. Using DUS, we release SOLAR 10.7B, an LLM with 10.7 billion parameters, that outperforms existing models like Llama 2 Touvron et al. (2023) and Mistral 7B Jiang et al. (2023) in various benchmarks. The proposed DUS method is immediately compatible with easy-to-use LLM frameworks such as HuggingFace Wolf et al. (2019) and does need additional changes to the training or inference framework. Furthermore, DUS is compatible with all transformer architectures, opening up new gateways to effectively and efficiently scale-up LLMs in a simple manner.
We have also developed SOLAR 10.7B-Instruct, a variant fine-tuned for tasks requiring strict adherence to complex instructions. It significantly outperforms the Mixtral-8x7B model across various evaluation metrics, evidencing an advanced proficiency that exceeds the capabilities of even larger models in terms of benchmark performance.
@@
2 SOLAR 10.7B Architectural Details
For developing LLMs with a better performance-to-size trade-off, we argue that the commonly used 7B-sized LLMs need to be at the Pareto-optimal curve and proceed to scale up the 7B LLMs. In doing so, we utilize pretrained weights of base models to scale up to larger LLMs in a more efficient manner. Specifically, we choose a well-performing base model and apply the novel DUS method to obtain a scaled-up model that utilizes the pretrained weights from the base model.
Note that other up-scaling methods, most notably MoE, require complex changes, such as a separate training framework optimized for MoE and custom CUDA kernels for efficient inference. In contrast, a model up-scaled with DUS can utilize the exact same training and inference framework as the base LLMs and still achieve maximal efficiency and efficacy. DUS offers a simple way to up-scale LLMs and any other transformer-based architectures, that is easy to use and compatible with existing widespread training and inference frameworks.
@@
Table 1: Training datasets used for the instruction and alignment tuning stages, respectively. For the instruction tuning process, we utilized the Alpaca-GPT4 Peng et al. (2023), OpenOrca Mukherjee et al. (2023), and Synth. Math-Instruct datasets, while for the alignment tuning, we employed the Orca DPO Pairs Intel (2023), Ultrafeedback Cleaned Ivison et al. (2023), and Synth. Math-Alignment datasets. The ‘Total # Samples‘ indicates the total number of samples in the entire dataset. The ‘Maximum # Samples Used‘ indicates the actual maximum number of samples that were used in training, which could be lower than the total number of samples in a given dataset. ‘Open Source‘ indicates whether the dataset is open-sourced.
@@
We have selected the 32-layer Llama 2 architecture as our base model, recognizing its robust and versatile framework as an optimal foundation for further advancements. We then initialize the Llama 2 architecture with pretrained weights from Mistral 7B, as it is one of the top performers compatible with the Llama 2 architecture.
@@
By adopting the Llama 2 architecture for our base model, we aim to leverage the vast pool of community resources while introducing novel modifications to further its capabilities.
Depth Up-Scaling.
@@
One naive way to up-scale the base LLM would be to repeat its layers once more, i.e., from $32$ to $64$ layers. This has the benefit that from layers 1 to 32 and from layers 33 to 64, there are no heterogeneity as those layers are taken directly from the base LLM. In other words, the ‘layer distance’, or the difference in the layer indices in the base model, is only bigger than 1 where layers 32 and 33 are connected, i.e., at the seam.
An additional benefit may be the potential for fast performance recovery after the up-scaling is done, as is also observed in MoE Komatsuzaki et al. (2022). The reason is that optimizing the up-scaled model could first focus on reducing the discrepancy in the layers at the seam, which would boost the performance of the up-scaled model rapidly. Then, a gradual optimization of all the layers would begin.
However, in the naive up-scaling approach, the layer distance at the seam reaches its maximum, potentially impeding the model’s ability to effectively utilize the pretrained weights. A potential solution to would be to sacrifice the middle layers, thereby reducing the discrepancy at the seam. Guided by this intuition, we devise the Depth Up-Scaling (DUS) method, which we illustrate in Figure 1.
In the first step of DUS, we take the base model, which is the 32-layer Llama2 architecture with Mistral 7B pretrained weights, and make a copy. Next, we slice off the last 8 layers from the original base model and the first 8 layers from the duplicate. This leaves us with two 24-layer models. In the final step, these models are concatenated to form a depth up-scaled model with 48 layers and 10.7 billion parameters.
The decision to remove 8 layers from each model was driven by our target performance-to-size trade-off. By discarding what would have been the middle layers in the up-scaled model, the layer distance at the seam is reduced as layer 24 of the first model to layer 9 of the second are connected instead of layer 32 and 1, respectively.
@@
Unlike the Mixture of Experts (MoE) approach, DUS does not require additional modules like a gating network or an expert selection process. Consequently, DUS models do not necessitate a distinct training framework for optimal training efficiency, nor do they require specialized CUDA kernels for fast inference. An LLM up-scaled with DUS can seamlessly integrate into existing training and inference frameworks while maintaining high efficiency. The following sections will delve into the training process of the up-scaled model, SOLAR 10.7B."
3 SOLAR 10.7B Training Details
Pretraining.
After DUS is applied to the base model, the performance initially drops below that of the base LLM. However, in-line with our hypothesis in Sec. 2, we observe fast performance recovery once we continually pretrain the up-scaled model. After continual pretraining, we perform fine-tuning in two stages: 1) instruction tuning and 2) alignment tuning.
@@
In the instruction tuning stage, the model is trained to follow instructions in a QA format Zhang et al. (2023b). We mostly use open-source datasets but also synthesize a math QA dataset to enhance the model’s mathematical capabilities. A rundown of how we crafted the dataset is as follows. First, seed math data are collected from the Math Hendrycks et al. (2021) dataset only, in order to avoid contamination with commonly used benchmark datasets such as GSM8K Cobbe et al. (2021). Then, using a process similar to MetaMath Yu et al. (2023), we rephrase the questions and answers of the seed math data. We use the resulting rephrased question-answer pairs as a QA dataset and call it ‘Synth. Math-Instruct‘.
@@
Model Merging.
Model merging is an effective way to boost model performance without further training. We merge some of the models that we trained in both the instruction and alignment tuning stages. We tried two model merging methods: 1) simple average of the weights and 2) SLERP Shoemake (1985).
4 Experimental Results
4.1 Training Datasets and Evaluation
@@
We present details regarding our training datasets for the instruction and alignment tuning stages in Tab.. 1. We do not always use the entirety of the dataset and instead subsample a set amount. We note that most of our training data is open-source, and the undisclosed datasets can be substituted for open-source alternatives such as the MetaMathQA Yu et al. (2023) dataset.
@@ 4.2 Main Results
We present evaluation results for our SOLAR 10.7B and SOLAR 10.7B-Instruct models along with other top-performing models in Tab. 2. SOLAR 10.7B outperforms other pretrained models of similar sizes, such as Qwen 14B and Mistral 7B, which shows that DUS is an effective method to up-scale base LLMs. Furthermore, despite the smaller size, SOLAR 10.7B-Instruct scores the highest in terms of H6, even surpassing the recent top-performing open-source LLM Mixtral 8x7B-Instruct-0.1 or Qwen 72B. The above results indicate DUS can up-scale models that are capable of achieving state-of-the-art performance when continually pretrained and fine-tuned. We also report data contamination results in Appendix C to show the integrity of SOLAR 10.7B-Instruct.
@@ 4.3.1 Instruction Tuning
We present ablation studies using different training datasets for the instruction tuning stage in Tab. 3. The ablated models are prefixed with SFT or supervised fine-tuning and are trained as follows. ‘SFT v1’ only uses the Alpaca-GPT4 dataset, whereas ‘SFT v2’ uses the OpenOrca dataset along with the Alpaca-GPT4 dataset. ‘SFT v3’ uses the Synth. Math-Instruct data when training along with the datasets used in ‘SFT v2’. Similarly, ‘SFT v4’ uses the Synth. Math-Instruct data when training along with the datasets used in ‘SFT v1’.
@@ 4.3.2 Alignment Tuning
This section describes our carefully designed alignment tuning strategy. We utilize Direct Preference Optimization (DPO) for practical alignment tuning. More specifically, we demonstrate the different training datasets used for training, the different SFT base models to initialize the DPO model, and finally, the model merging strategy to obtain the final alignment-tuned model.
@@
We ablate on the different alignment datasets used during DPO in Tab. 4. We use ‘SFT v3’ as the SFT base model for DPO. The description for the ablated models are as follows. ‘DPO v1’ only uses the Ultrafeedback Clean dataset while ‘DPO v2’ also used the Synth. Math-Alignment dataset.
@@ 5 Conclusion
This paper introduces depth up-scaling (DUS), a simple method to up-scale LLMs efficiently. Unlike other up-scaling techniques, such as MoE, DUS does not require specialized training or inference frameworks to achieve maximum efficiency. We demonstrate the effectiveness of DUS by training SOLAR 10.7B and its fine-tuned variant, SOLAR 10.7B-Instruct, which are depth up-scaled from smaller base LLMs. We hope the release of SOLAR 10.7B under the Apache 2.0 license can foster more collaborative research, improved accessibility, and sustainable growth in the field of NLP.
@@ Acknowledgements
We would like to extend our heartfelt gratitude to the teams at Hugging Face, particularly Clémentine Fourrier, Lewis Tunstall, Omar Sanseviero, and Philipp Schmid. Our appreciation also extends to the teams at AWS, notably Ritesh Vajaria, Gal Oshri, Jay Kwon, Brandon Lee, Effie Bae, and Rahul Sharma. We are grateful to the teams at Korea Telecom (KT), especially Jin Hyoung Lee, Jungsuk Park, Sungjoon Park, Hong-rae Wang, Kyeongsoo Jung, and Sunyoong Yoon, whose significant support has been instrumental in ensuring the broad compatibility of our model.
@@
Firstly, the model’s substantial computational requirements for training and inference may limit its accessibility for those with constrained computational resources. Secondly, despite efforts to reduce data contamination, SOLAR 10.7B, like all machine learning models, is susceptible to biases inherent in its training data, potentially leading to skewed outcomes in certain scenarios. Thirdly, the complexity of SOLAR 10.7B poses challenges in terms of interpretability and explainability, which can be problematic in applications where understanding the model’s decision-making process is critical.
Additionally, while SOLAR 10.7B performs admirably across various natural language processing tasks, its effectiveness can vary in different languages, particularly those with fewer resources, and in specialized domains. Moreover, the significant energy consumption required for training and running SOLAR 10.7B raises concerns about its environmental impact, a critical consideration in the context of sustainable AI development.
Finally, while SOLAR 10.7B-Instruct variant shows improved performance in following instructions, the model still necessitates task-specific fine-tuning for optimal performance in specialized applications, a process that can be both resource-intensive and not always effective. Acknowledging these limitations is essential for a comprehensive understanding of the proposed LLM’s capabilities and for guiding future research and development in LLMs.

### 29 Dec 2023 v2

@@ Abstract
We introduce SOLAR 10.7B, a large language model (LLM) with 10.7 billion parameters, demonstrating superior performance in various natural language processing (NLP) tasks. Inspired by recent efforts to efficiently up-scale LLMs, we present a method for scaling LLMs called depth up-scaling (DUS), which encompasses depthwise scaling and continued pretraining. In contrast to other LLM up-scaling methods that use mixture-of-experts, DUS does not require complex changes to train and inference efficiently. We show experimentally that DUS is simple yet effective in scaling up high-performance LLMs from small ones. Building on the DUS model, we additionally present SOLAR 10.7B-Instruct, a variant fine-tuned for instruction-following capabilities, surpassing Mixtral-8x7B-Instruct. SOLAR 10.7B is publicly available under the Apache 2.0 license, promoting broad access and application in the LLM field  [^1] https://huggingface.co/upstage/SOLAR-10.7B-v1.0.
@@ 1 Introduction
Figure 1: Depth up-scaling for the case with $n=32,s=48,$ and $m=8$. Depth up-scaling is achieved through a dual-stage process of depthwise scaling followed by continued pretraining.
The field of natural language processing (NLP) has been significantly transformed by the introduction of large language models (LLMs), which have enhanced our understanding and interaction with human language Zhang et al. (2023a). These advancements bring challenges such as the increased need to train ever larger models Rae et al. (2021); Wang et al. (2023); Pan et al. (2023); Lian (2023); Yao et al. (2023); Gesmundo and Maile (2023) owing to the performance scaling law Kaplan et al. (2020); Hernandez et al. (2021); Anil et al. (2023); Kaddour et al. (2023). To efficiently tackle the above, recent works in scaling language models such as a mixture of experts (MoE) Shazeer et al. (2017); Komatsuzaki et al. (2022) have been proposed. While those approaches are able to efficiently and effectively scale-up LLMs, they often require non-trivial changes to the training and inference framework Gale et al. (2023), which hinders widespread applicability. Effectively and efficiently scaling up LLMs whilst also retaining the simplicity for ease of use is an important problem Alberts et al. (2023); Fraiwan and Khasawneh (2023); Sallam et al. (2023); Bahrini et al. (2023).
Inspired by  Komatsuzaki et al. (2022), we present depth up-scaling (DUS), an effective and efficient method to up-scale LLMs whilst also remaining straightforward to use. DUS consists of scaling the base model along the depth dimension and continually pretraining the scaled model. Unlike Komatsuzaki et al. (2022), DUS does not scale the model using MoE and rather use a depthwise scaling method analogous to  Tan and Le (2019) which is adapted for the LLM architecture. Thus, there are no additional modules or dynamism as with MoE, making DUS immediately compatible with easy-to-use LLM frameworks such as HuggingFace Wolf et al. (2019) with no changes to the training or inference framework for maximal efficiency. Furthermore, DUS is applicable to all transformer architectures, opening up new gateways to effectively and efficiently scale-up LLMs in a simple manner. Using DUS, we release SOLAR 10.7B, an LLM with 10.7 billion parameters, that outperforms existing models like Llama 2 Touvron et al. (2023) and Mistral 7B Jiang et al. (2023) in various benchmarks.
We have also developed SOLAR 10.7B-Instruct, a variant fine-tuned for tasks requiring strict adherence to complex instructions. It significantly outperforms the Mixtral-8x7B-Instruct model across various evaluation metrics, evidencing an advanced proficiency that exceeds the capabilities of even larger models in terms of benchmark performance.
@@
2 Depth Up-Scaling
To efficiently scale-up LLMs, we aim to utilize pretrained weights of base models to scale up to larger LLMs Komatsuzaki et al. (2022). While existing methods such as Komatsuzaki et al. (2022) use MoE Shazeer et al. (2017) to scale-up the model architecture, we opt for a different depthwise scaling strategy inspired by Tan and Le (2019). We then continually pretrain the scaled model as just scaling the model without further pretraining degrades the performance.
@@ 2 Depth Up-Scaling
Table 1: Training datasets used for the instruction and alignment tuning stages, respectively. For the instruction tuning process, we utilized the Alpaca-GPT4 Peng et al. (2023), OpenOrca Mukherjee et al. (2023), and Synth. Math-Instruct datasets, while for the alignment tuning, we employed the Orca DPO Pairs Intel (2023), Ultrafeedback Cleaned Cui et al. (2023); Ivison et al. (2023), and Synth. Math-Alignment datasets. The ‘Total # Samples‘ indicates the total number of samples in the entire dataset. The ‘Maximum # Samples Used‘ indicates the actual maximum number of samples that were used in training, which could be lower than the total number of samples in a given dataset. ‘Open Source‘ indicates whether the dataset is open-sourced.
@@
Any $n$-layer transformer architecture can be used but we select the 32-layer Llama 2 architecture as our base model. We initialize the Llama 2 architecture with pretrained weights from Mistral 7B, as it is one of the top performers compatible with the Llama 2 architecture. By adopting the Llama 2 architecture for our base model, we aim to leverage the vast pool of community resources while introducing novel modifications to further enhance its capabilities.
Depthwise scaling.
From the base model with $n$ layers, we set the target layer count $s$ for the scaled model, which is largely dictated by the available hardware.
With the above, the depthwise scaling process is as follows. The base model with $n$ layers is duplicated for subsequent modification. Then, we remove the final $m$ layers from the original model and the initial $m$ layers from its duplicate, thus forming two distinct models with $n-m$ layers. These two models are concatenated to form a scaled model with $s=2\cdot(n-m)$ layers. Note that $n=32$ from our base model and we set $s=48$ considering our hardware constraints and the efficiency of the scaled model, i.e., fitting between 7 and 13 billion parameters. Naturally, this leads to the removal of $m=8$ layers. The depthwise scaling process with $n=32,s=48,$ and $m=8$ is depicted in ‘Step 1: Depthwise Scaling’ of Fig. 1.
We note that a method in the community that also scale the model in the same manner  [^2] https://huggingface.co/Undi95/Mistral-11B-v0.1 as ‘Step 1: Depthwise Scaling’ of Fig. 1 has been concurrently developed.
Continued pretraining.
The performance of the depthwise scaled model initially drops below that of the base LLM. Thus, we additionally apply the continued pretraining step as shown in ‘Step 2: Continued Pretraining’ of Fig. 1. Experimentally, we observe rapid performance recovery of the scaled model during continued pretraining, a phenomenon also observed in  Komatsuzaki et al. (2022). We consider that the particular way of depthwise scaling has isolated the heterogeneity in the scaled model which allowed for this fast performance recovery.
@@
Delving deeper into the heterogeneity of the scaled model, a simpler alternative to depthwise scaling could be to just repeat its layers once more, i.e., from $n$ to $2n$ layers. Then, the ‘layer distance’, or the difference in the layer indices in the base model, is only bigger than 1 where layers $n$ and $n+1$ are connected, i.e., at the seam.
However, this results in maximum layer distance at the seam, which may be too significant of a discrepancy for continued pretraining to quickly resolve. Instead, depthwise scaling sacrifices the $2m$ middle layers, thereby reducing the discrepancy at the seam and making it easier for continued pretraining to quickly recover performance. We attribute the success of DUS to reducing such discrepancies in both the depthwise scaling and the continued pretraining steps. We also hypothesize that other methods of depthwise scaling could also work for DUS, as long as the discrepancy in the scaled model is sufficiently contained before the continued pretraining step.
Comparison to other up-scaling methods.
@@
Unlike  Komatsuzaki et al. (2022), depthwise scaled models do not require additional modules like gating networks or dynamic expert selection. Consequently, scaled models in DUS do not necessitate a distinct training framework for optimal training efficiency, nor do they require specialized CUDA kernels for fast inference. A DUS model can seamlessly integrate into existing training and inference frameworks while maintaining high efficiency.
3 Training Details
After DUS, including continued pretraining, we perform fine-tuning of SOLAR 10.7B in two stages: 1) instruction tuning and 2) alignment tuning.
@@ 3 Training Details
In the instruction tuning stage, the model is trained to follow instructions in a QA format Zhang et al. (2023b). We mostly use open-source datasets but also synthesize a math QA dataset to enhance the model’s mathematical capabilities. A rundown of how we crafted the dataset is as follows. First, seed math data are collected from the Math Hendrycks et al. (2021) dataset only, to avoid contamination with commonly used benchmark datasets such as GSM8K Cobbe et al. (2021). Then, using a process similar to MetaMath Yu et al. (2023), we rephrase the questions and answers of the seed math data. We use the resulting rephrased question-answer pairs as a QA dataset and call it ‘Synth. Math-Instruct‘.
@@
4 Results
4.1 Experimental Details
@@
We present details regarding our training datasets for the instruction and alignment tuning stages in Tab. 1. We do not always use the entire dataset and instead subsample a set amount. Note that most of our training data is open-source, and the undisclosed datasets can be substituted for open-source alternatives such as the MetaMathQA Yu et al. (2023) dataset.
@@
Model merging.
Model merging methods such as  Yadav et al. (2023) can boost model performance without further training. We merge some of the models that we trained in both the instruction and alignment tuning stages. We implement our own merging methods although popular open source also exist such as MergeKit [^3] https://github.com/cg123/mergekit.
@@ 4.2 Main Results
We present evaluation results for our SOLAR 10.7B and SOLAR 10.7B-Instruct models along with other top-performing models in Tab. 2. SOLAR 10.7B outperforms other pretrained models of similar sizes, such as Qwen 14B and Mistral 7B, which shows that DUS is an effective method to up-scale base LLMs. Furthermore, despite the smaller size, SOLAR 10.7B-Instruct scores the highest in terms of H6, even surpassing the recent top-performing open-source LLM Mixtral 8x7B-Instruct-v0.1 or Qwen 72B. The above results indicate DUS can up-scale models that are capable of achieving state-of-the-art performance when fine-tuned. We also report data contamination results for SOLAR 10.7B-Instruct in Appendix C.
@@ 4.3.1 Instruction Tuning
We present ablation studies using different training datasets for the instruction tuning in Tab. 3. The ablated models are prefixed with SFT for supervised fine-tuning. ‘SFT v1’ only uses the Alpaca-GPT4 dataset, whereas ‘SFT v2’ also uses the OpenOrca dataset. ‘SFT v3’ uses the Synth. Math-Instruct dataset along with the datasets used in ‘SFT v2’. Similarly, ‘SFT v4’ uses the Synth. Math-Instruct dataset along with the datasets used in ‘SFT v1’.
@@ 4.3.2 Alignment Tuning
As we utilize DPO for practical alignment tuning, there are additional aspects to ablate such as the SFT base models used. Thus, we present ablations for the different training datasets used for training, the different SFT base models to initialize the DPO model, and finally, the model merging strategy to obtain the final alignment-tuned model.
@@
We ablate on the different alignment datasets used during DPO in Tab. 4. We use ‘SFT v3’ as the SFT base model for DPO. ‘DPO v1’ only uses the Ultrafeedback Clean dataset while ‘DPO v2’ also used the Synth. Math-Alignment dataset.
@@ 5 Conclusion
We introduce SOLAR 10.7B and its fine-tuned variant SOLAR 10.7B-Instruct, which are depth up-scaled (DUS) models with 10.7 billion parameters. They show superior performance over models like Llama 2, Mistral 7B, and Mixtral-7B-Instruct in essential NLP tasks while maintaining computational efficiency. Thus, DUS is effective in scaling-up highly performant LLMs from smaller ones. With more exploration, DUS could be further improved, paving a new path to efficiently scaling LLMs.
@@ Acknowledgements
We would like to extend our gratitude to the teams at Hugging Face, particularly Clémentine Fourrier, Lewis Tunstall, Omar Sanseviero, and Philipp Schmid. Our appreciation also extends to the teams at AWS, notably Ritesh Vajaria, Gal Oshri, Jay Kwon, Brandon Lee, Effie Bae, and Rahul Sharma. We are grateful to the teams at Korea Telecom (KT), especially Jin Hyoung Lee, Jungsuk Park, Sungjoon Park, Hong-rae Wang, Kyeongsoo Jung, and Sunyoong Yoon, whose significant support has been instrumental in ensuring the broad compatibility of our model. Additionally, we would like to extend our thanks to the open community for their invaluable contributions and feedback.
@@
Our study on the Depth Up-Scaling (DUS) has important limitations and considerations. One key limitation is the need for more thorough explorations of hyperparameters used in the DUS approach. Namely, we removed $m=8$ layers from both ends of our base model, primarily due to hardware limitations. However, we have not yet determined if this value is optimal for enhancing performance. The extended time and cost of continued pretraining made it challenging to conduct more comprehensive experiments, which we aim to address in future work through various comparative analyses.
In terms of the model’s broader implications, there are several points to note. The model’s significant computational demands for training and inference might limit its use, especially for those with restricted computational resources. Additionally, like all machine learning models, it is vulnerable to biases in its training data, which could lead to skewed outcomes in certain situations. Furthermore, the substantial energy consumption required for training and operating the model raises environmental concerns, which are critical in the pursuit of sustainable AI development.
Lastly, while the fine-tuned variant of the model shows improved performance in following instructions, it still requires task-specific fine-tuning for optimal performance in specialized applications. This fine-tuning process can be resource-intensive and not always effective. Recognizing and addressing these limitations is essential for a comprehensive understanding of the proposed Large Language Model’s capabilities and for guiding future research and development in the field of LLMs.